NvidiaIntely (nvidiaintely):

Strontium-90 is a radioisotope that will decrease in mass by one-half every 29 years. How many years will it take for a 10.0-gram sample of strontium-90 to decay to 5.00 grams? 1 14.5 29 58

NvidiaIntely (nvidiaintely):

@Kevin @agent0smith @welshfella @Will.H @sweetburger @thephysicsman @TheSmartOne

OpenStudy (welshfella):

5 g is half of 10 g so......

OpenStudy (kevin):

5/10 = (1/2)^t/29 t = 29

NvidiaIntely (nvidiaintely):

Thanks Welshfella and Kevin! (That was good hint Welshfella)

OpenStudy (welshfella):

yw